puppylover1234:

Sam can collect 30 cups of water in 5 minutes. If he continues at this rate, how many cups of water could Sam collect in 12 minutes? Group of answer choices A)60 B)72 C)360 D)12

AZ:

We can set up a proportion for this since he's doing it at the same rate 30 cups in 5 minuets ?? cups in 12 minutes \(\dfrac{30}{5} = \dfrac{x}{12}\) do you know how to cross multiply and solve for x?

puppylover1234:

i think the answer is 72

AZ:

That's correct!

puppylover1234:

oh thank you!

AZ:

so remember cross multiplying is like this \(\dfrac{a}{b} = \dfrac{c}{d}\) we get \( ad = bc\) so in your question we would have gotten 30 * 12 = 5 * x and then when you solve for x, you would get the answer that you got!

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @puppylover1234 oh thank you! \(\color{#0cbb34}{\text{End of Quote}}\) It was my pleasure!